OpenStudy (anonymous):

how to answer this? 5(2^(3x-1))-7=13

OpenStudy (anonymous):

5(2^(3x-1))=20 (2^(3x-1))=4 (2^(3x-1))=2^2 3x - 1 = 2 3x = 3 x = 1

OpenStudy (anonymous):

\[5(2^{3x-1})\]